Mpohlo bought nine shirts at a total cost of R2 520,00. The large shirts costs R300,00 each and the small


shirts costs R240,00 each. Choose the correct statement.


[1] He bought twice as many small shirts as large shirts


[2] He bought three large shirts


[3] He paid R720 in total for the small shirts


[4] He paid double the amount for the large shirts as the amount that he paid for the small shirts


1
Expert's answer
2022-04-17T23:01:45-0400

Let x - amount of small shirts, then amount of large shirts is 9 - x

"x\cdot240+(9-x)\cdot300=2520\\
2700-60x=2520\\
x=\frac{2700-2520}{60}=\frac{180}{60}=3"

Amount of small shirts is 3, amount of large shirts is 9 - 3 = 6.

The correct statement is 3, that total cost of small shirts is "3\cdot240=720"

Answer: [3] He paid R720 in total for the small shirts.